OpenStudy (anonymous):

can someone show me why 1 statcoulomb=3.3356*10^(-10) C

OpenStudy (anonymous):

In cgs units coulomb force is measured in dynes and charge in statcoulombs. In mks units coulomb force is measured in newtons and charge in coulombs. 10^5 dynes = 1 newton writing coulombs law \[ F =10^{5 }*8.987 *10^{9} Q _{c}^{2}/D ^{2}(m) \] writing out all the dimensions in mks we get Nt*m^2/C^2 In cgs unit F= q^2/d^2 dynes*cm^2/statcoul^2 set F in cgs = F in mks q^2/d^2 = 8.987*10^9*Q^2/D^2 Thus equating the units Dynes cm^2/statcoul^2= 8.987*10^9*NT*m^2/coul^2 let k=coul/statcoul ;100cm =1m: 1Nt=10^5dynes Dynes*.0001m^2/statcoul^2 = 8.897*10^9*10^5*dynes*m^2/(k*statcoul)^2 thus k^2 = 8.987*10^9*10^9 and k = 2.9827*10^9 Thus q = Q/k = 3.3356*10^-10

OpenStudy (anonymous):

You can't imagine how many times I stumbled in this response.